# Thermally activated particle motion in biased correlated Gaussian disorder potentials

### v2

@@ Abstract
Thermally activated particle motion in disorder potentials is controlled by the large-$\Delta V$ tail of the distribution of height $\Delta V$ of the potential barriers created by the disorder. We employ the optimal fluctuation method to evaluate this tail for correlated quenched Gaussian potentials in one dimension in the presence of a small bias of the potential. We focus on the mean escape time (MET) of overdamped particles averaged over the disorder. We show that the bias leads to a strong (exponential) reduction of the MET in the direction along the bias. The reduction depends both on the bias, and on detailed properties of the covariance of the disorder, such as its derivatives and asymptotic behavior at large distances. We verify our theoretical predictions, as well as earlier predictions for zero bias, by performing large-deviation simulations of the potential disorder. The simulations employ correlated random potential sampling based on the circulant embedding method and the Wang-Landau algorithm, which enable us to probe probability densities smaller than $10^{-1200}$.
@@ I Introduction
In order to verify our theoretical predictions, as well as the earlier predictions LV ; M2022 for zero bias, we perform large-deviation simulations of the potential disorder. The simulations employ correlated random potential sampling based on the circulant embedding method and the Wang-Landau algorithm, which enable us to probe probability densities smaller than $10^{-1200}$. As we will show, the simulation results strongly support the theory.
@@ II Optimal Fluctuation Method
Assuming that the optimal configuration of $V\left(x\right)$ is smooth, we can write down the conditions specifying the potential barrier $\Delta V$ on an interval of length $2L$:
@@ II.2.1 Monotonically decreasing covariance
Let us first examine how the presence of a small bias affects the optimal value of $L=L(E)$ as described by Eq. (21). One can see from Eq. (21) that, as $E$ goes to zero, $\kappa^{\prime}(L)$ also goes to zero, so that the optimal barrier width $2L$ goes to infinity LV ; M2022 . However, as one can check a posteriori, it does so slower than $1/E$, that is $\lim_{E\rightarrow 0^{+}}EL\left(E\right)=0$, and we will rely on this property.
@@
The non-analytic correction $\sim E^{\frac{\alpha}{\alpha+1}}$, coming from the bias, describes an increase in the “action cost” of creating the barrier $\Delta V$ and, as a result, a decrease in the probability $\mathcal{P}(\Delta V)$ of observing this barrier. This perturbative calculation demands the strong inequality $E\ll\Delta V$. When $E\rightarrow 0$, the first line of Eq.  (26) is reproduced, as to be expected.
@@ III.1 Simulation method
Instead of the poorly scalable matrix multiplications, a more efficient way to sample HGFs is to use the Circulant Embedding method (CEM) Chan1994 ; Dietrich1997 . The method is based on embedding a covariance matrix $C_{i,j}$ of size $M\times M$ in a larger circulant covariance matrix $C^{\prime}_{ij}$ of size $M^{\prime}>2M$. Then, using the fact that a matrix multiplication with a circulant matrix implements convolution, one can replace the matrix equation (48) with multiplication in Fourier space. Therefore the required data (like the result of matrix diagonalization) can be easily pre-computed numerically by using the fast Fourier transform (FFT), while sampling a HGF requires only $\mathcal{O}(M^{\prime}\ln M^{\prime})$ operations. Some instructive implementations of the CEM for large deviation simulations of fractal Brownian motion (whose time derivative is a stationary Gaussian process) can be found in Refs. Hartmann_2 ; Hartmann_1 .
Here we are interested in extremely small probability densities, which are virtually impossible to reach with conventional Monte Carlo (MC) simulations employing the Metropolis-Hastings algorithm. Therefore, in order to reach the large-$\Delta V$ tail of the distribution $\mathcal{P}(\Delta V)$, we employed the Wang-Landau (WL) algorithm WL1 ; WL2 . Unlike the ordinary Metropolis-Hastings algorithm, where the acceptance/rejection decisions are Markovian, the Wang-Landau algorithm takes into account information about previously visited states in such a way that it “forces” the algorithm to explore the available configurational space more quickly.
In a nutshell, the WL algorithm aims at estimating the density of states $\mathcal{P}(\Delta V)=\exp[-S(\Delta V)]$ on a chosen interval $a\leq\Delta V\leq b$, updating on each MC step the histogram of visited states $H(\Delta V)$ and adjusting the entropy function $S(\Delta V)$ in an iterative manner, see Ref. Landau_Binder for details.
At the start of the simulation, the histogram is initialized to zero, $H(\Delta V)=0$, and the entropy is set to some guess function (we use $S(\Delta V)=1$). Let $\bigl{\{}\vec{\xi}_{\text{r}},\vec{V}_{\text{r}},\Delta V_{\text{r}}\bigl{\}}$ represent the running configurations of the random vector, the disorder potential computed using the CEM, and the maximal potential barrier of $\vec{V}_{\text{r}}$, respectively. A proposed configuration of the random vector $\vec{\xi}_{\text{p}}$ is generated by changing a randomly chosen component $[\vec{\xi}_{\text{r}}]_{j}$ of the random vector according to the Gaussian distribution centered at $[\vec{\xi}_{\text{r}}]_{j}$:
@@
where $\overline{H(\Delta V)}$ is the mean value of the histogram. Once this condition is met, $f$ is reduced, $f\to f/2$, and the histogram is reset to zero. Then, the process continues until $f$ is sufficiently small.
It is known that, at early stages of a simulation, the Wang-Landau algorithm violates the detailed balance condition. Asymptotically, however, the detailed balance is recovered as the modification factor $f$ tends to zero Landau_Binder . We stopped the Wang-Landau simulations after 18 reductions of the modification factor $f$, which suffices for our purposes of large deviation simulations. This is because the statistical errors of the Wang-Landau algorithm, known to saturate Zhou , are of the order $\mathcal{O}(1)$, which is much smaller than the resulting entropy in our simulations, $S(\Delta V)\sim\mathcal{O}(10^{3})$.
@@ III.2 Simulation results
To verify our theoretical predictions for the monotonic and nonmonotonic covariances, we implemented the Wang-Landau algorithm of sampling discretized configurations of the disorder Gaussian potential $V(x)$ on a regular lattice of length $M=10^{3}$ with the following three covariances:
@@ III.2.1 Zero bias
Examples of configurations of disorder potential $V(x)$, conditioned on large $\Delta V$ and sampled in the WL simulations, are presented in Fig. 4. Also shown are the optimal configurations predicted by Eq. (24) with the optimal values of $L$. As one can see, the agreement of theory and simulations is excellent in all the three cases. Notice that, for the NMN and NMP covariances, the theoretically predicted optimal barrier size is finite, allowing the true minimum of the action to be reached in the simulations. In the case of MD covariance, the true minimum of the action can be reached only when $L\to\infty$, and it is therefore inaccessible in numerical simulations. A finite value of $L$, observed in the simulations, is caused by finite-size effects as it is comparable with the size of the simulated system.
@@ IV Summary and Discussion
We found that the presence of a small potential bias leads to an exponentially large reduction in the MET of overdamped particles trapped in local potential minima. The leading-order correction, which describes this reduction, behaves differently in disorder potentials with monotonic and non-monotonic covariance.
In the non-monotonic case, the effect of the bias can be accounted for via a perturbative expansion in the bias. In the monotonic case, the scaling of the MET with the bias is nontrivial, as it is affected by the large-distance asymptotic of the inverse function of the derivative of the covariance. The optimal barrier width of the biased potential in this case is finite, in contrast to the unbiased case, where it is infinite. Even a very small potential bias has a strong effect on the characteristic barrier width. As a result, all bias-related effects are more pronounced for disorder potentials with monotonically decreasing covariances.
We verified our predictions, as well as earlier predictions LV ; M2022 for zero bias, in numerical simulations. The simulations employed the Wang-Landau algorithm and the circulant embedding method of sampling a homogeneous Gaussian field. We measured the large-$\Delta V$ tail of the barrier distribution $\mathcal{P}\left(\Delta V\right)$ for different covariances and bias magnitudes. The method also allowed us to sample the disorder potentials $V(x)$, allowing for a direct comparison with the OFM predictions for the optimal configurations, demonstrating excellent agreement. Combining the Wang-Landau algorithm with the circulant embedding method, we were able to measure probability densities below $10^{-1200}$. The numerical methods, which we employed here, should be also useful when studying large deviation statistics of other Gaussian processes and fields.

### v3

@@ Abstract
Thermally activated particle motion in disorder potentials is controlled by the large-$\Delta V$ tail of the distribution of height $\Delta V$ of the potential barriers created by the disorder. We employ the optimal fluctuation method to evaluate this tail for correlated quenched Gaussian potentials in one dimension in the presence of a small bias of the potential. We focus on the mean escape time (MET) of overdamped particles averaged over the disorder. We show that the bias leads to a strong (exponential) reduction of the MET in the direction along the bias. The reduction depends both on the bias, and on detailed properties of the covariance of the disorder, such as its derivatives and asymptotic behavior at large distances. We verify our theoretical predictions for the large-$\Delta V$ tail of the barrier height distribution, as well as earlier predictions of this tail for zero bias, by performing large-deviation simulations of the potential disorder. The simulations employ correlated random potential sampling based on the circulant embedding method and the Wang-Landau algorithm, which enable us to probe probability densities smaller than $10^{-1200}$.
@@ I Introduction
To verify our theoretical predictions for the large-$\Delta V$ tail of the barrier height distribution, as well as the earlier predictions of this tail for zero bias  LV ; M2022 , we perform large-deviation simulations of the potential disorder. The simulations employ correlated random potential sampling based on the circulant embedding method and the Wang-Landau algorithm, which enable us to probe probability densities smaller than $10^{-1200}$. As we will show, the simulation results strongly support the theory.
@@ II Optimal Fluctuation Method
Assuming that the optimal configuration of $V\left(x\right)$ is smooth, we can write down the conditions specifying the potential barrier $\Delta V$ on an interval $|x|<L$ of an a priori unknown length $2L$ arbitraryconstant :
@@ II.2.1 Monotonically decreasing covariance
Let us first examine how the presence of a small bias affects the optimal value of $L=L(E)$ as described by Eq. (21). One can see from Eq. (21) that, as $E$ goes to zero, $\kappa^{\prime}(2L)$ also goes to zero, so that the optimal barrier width $2L$ goes to infinity LV ; M2022 . However, as one can check a posteriori, it does so slower than $1/E$, that is $\lim_{E\rightarrow 0^{+}}EL\left(E\right)=0$, and we will rely on this property.
@@
The nonanalytic correction $\sim E^{\frac{\alpha}{\alpha+1}}$, coming from the bias, describes an increase in the “action cost” of creating the barrier $\Delta V$ and, as a result, a decrease in the probability $\mathcal{P}(\Delta V)$ of observing this barrier. This perturbative calculation demands the strong inequality $E\ll\Delta V$. When $E\rightarrow 0$, the first line of Eq.  (26) is reproduced, as to be expected.
@@ III.1 Simulation method
Instead of the poorly scalable matrix multiplications, a more efficient way to sample HGFs is to use the Circulant Embedding method (CEM) Chan1994 ; Dietrich1997 . The method is based on embedding a covariance matrix $C_{i,j}$ of size $M\times M$ in a larger circulant covariance matrix $C^{\prime}_{ij}$ of size $M^{\prime}>2M$. Then, using the fact that a matrix multiplication with a circulant matrix implements convolution, one can replace the matrix equation (48) with multiplication in Fourier space. Therefore the required data (like the result of matrix diagonalization) can be easily pre-computed numerically by using the fast Fourier transform (FFT), while sampling a HGF requires only $\mathcal{O}(M^{\prime}\ln M^{\prime})$ operations. In our implementation of the CEM we allowed for free ends of the sampled HGF $V(x)$. One could instead use periodic boundary conditions $V(x+M)=V(x)$ periodic . See also Refs. Hartmann_2 ; Hartmann_1 for instructive implementations of the CEM for large deviation simulations of fractal Brownian motion (whose time derivative is a stationary Gaussian process).
Here we are interested in extremely small probability densities, which are virtually impossible to reach with conventional Monte Carlo (MC) simulations employing the Metropolis-Hastings algorithm. Therefore, in order to reach the large-$\Delta V$ tail of the distribution $\mathcal{P}(\Delta V)$, we employed the Wang-Landau (WL) algorithm WL1 ; WL2 . Unlike the ordinary Metropolis-Hastings algorithm, where the acceptance/rejection decisions are Markovian, the WL algorithm takes into account information about previously visited states in such a way that it “forces” the algorithm to explore the available configurational space more quickly.
In a nutshell, the WL algorithm aims at estimating the density of states $\mathcal{P}(\Delta V)=\exp[-S(\Delta V)]$ on a chosen interval $a\leq\Delta V\leq b$, updating on each MC step the histogram of visited states $H(\Delta V)$ and adjusting the action $S(\Delta V)$ in an iterative manner, see Ref. Landau_Binder for details.
At the start of the simulation, the histogram is initialized to zero, $H(\Delta V)=0$, and the action is set to some guess function (we use $S(\Delta V)=1$). Let $\bigl{\{}\vec{\xi}_{\text{r}},\vec{V}_{\text{r}},\Delta V_{\text{r}}\bigl{\}}$ represent the running configurations of the random vector, the disorder potential computed using the CEM, and the maximal potential barrier of $\vec{V}_{\text{r}}$, respectively. A proposed configuration of the random vector $\vec{\xi}_{\text{p}}$ is generated by changing a randomly chosen component $[\vec{\xi}_{\text{r}}]_{j}$ of the random vector according to the Gaussian distribution centered at $[\vec{\xi}_{\text{r}}]_{j}$:
@@
where $\overline{H(\Delta V)}$ is the mean value of the histogram flat_note . Once this condition is met, $f$ is reduced, $f\to f/2$, and the histogram is reset to zero. Then, the process continues until $f$ is sufficiently small.
It is known that, at early stages of a simulation, the WL algorithm violates the detailed balance condition. The detailed balance is recovered, however, as the modification factor $f$ tends to zero Landau_Binder . We stopped the simulations after 18 reductions of the modification factor $f$. The resulting relative accuracy of the simulations, $1-S^{(17)}(\Delta V)/S^{(18)}(\Delta V)$ was of the order of $10^{-3}$, where $S^{(i)}(\Delta V)$ is the numerical estimate of the action corresponding to the $i$-th reduction of the modification factor $f$. The remaining statistical errors of the WL algorithm, which are known to saturate at a nonzero value (which depends on the protocol of reduction $f$ Zhou ; Belardinelli ), were of order $\mathcal{O}(1)$. In its turn, the numerically estimated action $S(\Delta V)$ varied from hundreds to thousands. The resulting relative accuracy was sufficient for our purposes.
@@ III.2 Simulation results
To verify our theoretical predictions for the tail of the barrier height distribution and the corresponding optimal configurations of the disorder for the monotonic and nonmonotonic covariances, we implemented the WL algorithm of sampling discretized configurations of the disorder Gaussian potential $V(x)$ on a regular lattice of length $M=10^{3}$ with the following three covariances:
@@ III.2.1 Zero bias
Examples of configurations of disorder potential $V(x)$, conditioned on large $\Delta V$ and sampled in the WL simulations shifts , are presented in Fig. 4. Also shown are the optimal configurations predicted by Eq. (24) with the optimal values of $L$. As one can see, the agreement of theory and simulations is excellent in all the three cases. Notice that, for the NMN and NMP covariances, the theoretically predicted optimal barrier size is finite, allowing the true minimum of the action to be reached in the simulations. In the case of MD covariance, the true minimum of the action can be reached only when $L\to\infty$, and it is therefore inaccessible in numerical simulations. A finite value of $L$, observed in the simulations, is caused by finite-size effects as it is comparable with the size of the simulated system.
@@ IV Summary and Discussion
We found that the presence of a small potential bias leads to an exponentially large reduction in the MET of overdamped particles trapped in local potential minima. The leading-order correction, which describes this reduction, behaves differently in disorder potentials with monotonic and nonmonotonic covariance.
In the nonmonotonic case, the effect of the bias can be accounted for via a perturbative expansion in the bias. In the monotonic case, the scaling of the MET with the bias is nontrivial, as it is affected by the large-distance asymptotic of the inverse function of the derivative of the covariance. The optimal barrier width of the biased potential in this case is finite, in contrast to the unbiased case, where it is infinite. Even a very small potential bias has a strong effect on the characteristic barrier width. As a result, all bias-related effects are more pronounced for disorder potentials with monotonically decreasing covariances.
We verified in numerical simulations our predictions for the large-$\Delta V$ tail of the barrier height distribution, as well as earlier predictions of this tail for zero bias LV ; M2022 . The simulations employed the WL algorithm and the circulant embedding method of sampling a homogeneous Gaussian field. We measured the large-$\Delta V$ tail of the barrier distribution $\mathcal{P}\left(\Delta V\right)$ for different covariances and bias magnitudes. The method also allowed us to sample the disorder potentials $V(x)$, allowing for a direct comparison with the OFM predictions for the optimal configurations, demonstrating excellent agreement. Combining the WL algorithm with the circulant embedding method, we were able to measure probability densities below $10^{-1200}$. The numerical methods, which we employed here, should be also useful when studying large deviation statistics of other Gaussian processes and fields.
@@
Finally, it is worth reminding that our results for the distribution tail of $\mathcal{P}(\Delta V)$ are readily applicable to the distribution tail of the differences between adjacent maxima and minima of correlated Gaussian processes in time.